Write the equilibrium-constant expression for the reaction shown in terms of
[A],[B],[C],and [D].
3A(g)+B(g)↽−−⇀2C(g)+2D(g)
1
Expert's answer
2019-08-02T05:45:56-0400

For the given reaction

"3A_{(g)}+B_{(g)} \rightleftharpoons 2C_{(g)} + 2D_{(g)}"

equilibrium-constant expression will be

"K_c=\frac{[C]^2 \cdot [D]^2}{[A]^3 \cdot [B]}"

where [A], [B], [C], and [D] are molar concentrations of A, B, C, and D respectively.